OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x − 3y = 3 5x − 4y = 3 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system −3 3

OpenStudy (anonymous):

1) you want to get rid of the y's 2) 1st you multiply to first equation by 4 to get 16x-12y=3 3) then you multiply the second equation by -3 to get -15x+12y=-9 4) add both of the equations and automatically the y's cancel out 5) you are left with x=-6 Hope that helps :D

OpenStudy (amistre64):

your "2)" is off i bit; 4*3 = 12 on the end, not 3

OpenStudy (anonymous):

thnx .... i was about to change that :)

OpenStudy (amistre64):

the concepts you present are good, the mathing needs work tho ;) so #5 would need to be adjusted as well